# Existence of solution to a system of PDEs modeling the crystal growth inside lithium batteries

### v2

@@ 3.2. Lemma (stability of the time-discrete order parameter)
$(u^{k},(u^{k})^{-})+\tau(\boldsymbol{{A}}(\nabla u^{k})\nabla u^{k},\nabla(u^{k})^{-})-\tau((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),(u^{k})^{-})\\ -\tau((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),(u^{k})^{-})+\tau g^{\prime}(u^{k}),(u^{k})^{-})=(u^{k-1},(u^{k})^{-})$ (3.6)
@@
$\|(u^{k})^{-}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\tau c_{A}|(u^{k})^{-}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq(u^{k-1},(u^{k})^{-})+\tau((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),(u^{k})^{-})\\ +\tau((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),(u^{k})^{-})-\tau(g^{\prime}(u^{k}),(u^{k})^{-}).$ (3.7)
By assumption $(u^{k-1})^{-}=0$ and noting from (2.12) and (3.2) that $(g^{\prime}(u^{k}),(u^{k})^{-})\geq 0$, $((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),(u^{k})^{-})\leq 0$ and $((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),(u^{k})^{-})\leq 0$ it follows that
$\|(u^{k})^{-}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq 0\leavevmode\nobreak\ \leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ u^{k}\geq 0.$ (3.8)
@@ 3.3. Remark (useful inequalities)
${\left|\left((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k})+(m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),\chi\right)\right|}\leq C\|u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\chi\|_{\operatorname{L}_{2}(\Omega)}.$ (3.13)
3.4. Lemma (existence and uniqueness of solutions for the time-discrete order parameter)
@@
$(d_{\tau}u^{k},\chi)+I^{\prime}(u^{k})\chi=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\chi\in\operatorname{H}^{1}(\Omega),$ (3.14)
@@
$I^{\prime}(w)\chi=\left(\boldsymbol{{A}}(\nabla w)\nabla w,\nabla\chi\right)+\left(g^{\prime}(w)-(m(c^{k-1}))^{-}h_{1}^{\prime}(w)-(m(c^{k-1}))^{+}h_{2}^{\prime}(w),\chi\right)$ (3.15)
is the (Fréchet) derivative of the energy functional $I:\operatorname{H}^{1}(\Omega)\to\mathbb{R}$ with
$I(w)=\int_{\Omega}\Big{(}\frac{a(\nabla w)^{2}}{2}|\nabla w|^{2}+g^{\prime}(w)-(m(c^{k-1}))^{-}h_{1}^{\prime}(w)-(m(c^{k-1}))^{+}h_{2}^{\prime}(w)\Big{)}.$ (3.16)
@@
$I_{k}(w)=\frac{1}{2\tau}\|w-u^{k-1}\|^{2}_{\operatorname{L}_{2}(\Omega)}+I(w),$ (3.17)
which is the functional from which we derive (3.14). From Theorem 3.30 in Dacorogna (2008), to show the existence of a minimiser of (3.17) it is enough to prove that the function $F:\Omega\times\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$, defined by
$F(x,\xi,\boldsymbol{p}):=\frac{a(\boldsymbol{p})^{2}}{2}|\boldsymbol{p}|^{2}+\frac{1}{2\tau}(\xi-u^{k-1})^{2}\\ +g^{\prime}(\xi)-(m(c^{k-1}))^{-}h_{1}^{\prime}(\xi)-(m(c^{k-1}))^{+}h_{2}^{\prime}(\xi),$ (3.18)
is a Carathéodory function, coercive and satisfies a sufficient growth condition. From (Bartels, 2015, Remark 2.4(i)) it follows that $F$ is a Carathédory function. From (3.11) we deduce that there are constants $\gamma_{1},\gamma_{2}>0$ such that
$-\gamma_{1}\leq g^{\prime}(\xi)-(m(c^{k-1}))^{-}h_{1}^{\prime}(\xi)-(m(c^{k-1}))^{+}h_{2}^{\prime}(\xi)\leq\gamma_{2}.$ (3.19)
@@
$F(x,\xi,\boldsymbol{p})$ $=\frac{a(\boldsymbol{p})^{2}}{2}|\boldsymbol{p}|^{2}+\frac{1}{2\tau}(\xi-u^{k-1})^{2}+g^{\prime}(\xi)-(m(c^{k-1}))^{-}h_{1}^{\prime}(\xi)$
$\qquad-(m(c^{k-1}))^{+}h_{2}^{\prime}(\xi)$
@@
$F(x,\xi,\boldsymbol{p})$ $=\frac{a(\boldsymbol{p})^{2}}{2}|\boldsymbol{p}|^{2}+\frac{1}{2\tau}(\xi-u^{k-1})^{2}+g^{\prime}(\xi)-(m(c^{k-1}))^{-}h_{1}^{\prime}(\xi)$
$\qquad-(m(c^{k-1}))^{+}h_{2}^{\prime}(\xi)$
@@
$\int_{\Omega}\frac{1}{\tau}(w_{1}-w_{2})v+\int_{\Omega}(\boldsymbol{{A}}(\nabla w_{1})\nabla w_{1}-\boldsymbol{{A}}(\nabla w_{2})\nabla w_{2})\cdot\nabla v\\ +\,\int_{\Omega}\Big{(}g^{\prime}(w_{1})-g^{\prime}(w_{2})\Big{)}v-\int_{\Omega}(m(c^{k-1}))^{+}\Big{(}h_{1}^{\prime}(w_{1})-h_{1}^{\prime}(w_{2})\Big{)}v\\ -\int_{\Omega}(m(c^{k-1}))^{-}\Big{(}h_{2}^{\prime}(w_{1})-h_{2}^{\prime}(w_{2})\Big{)}v=0.$ (3.22)
Choosing $v=w_{1}-w_{2}$ in (3.22), using the Lipschitz continuity of $g^{\prime}$, $h_{1}^{\prime}(u^{k})$ and $h_{2}^{\prime}(u^{k})$, (3.11), and noting from Remark 2.6 the monotonicity of the mapping $w\mapsto J_{a}^{\prime}(w)$, we obtain
$\frac{1}{\tau}\|w_{1}-w_{2}$ $\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq\int_{\Omega}|(g^{\prime}(w_{1})-g^{\prime}(w_{2}))(w_{1}-w_{2})|$ (3.23)
$+C\int_{\Omega}|h_{1}^{\prime}(w_{1})-h_{1}^{\prime}(w_{2})||w_{1}-w_{2}|\leq C^{\star}\|w_{1}-w_{2}\|^{2}_{\operatorname{L}_{2}(\Omega)}$
$+C\int_{\Omega}|h_{2}^{\prime}(w_{1})-h_{2}^{\prime}(w_{2})||w_{1}-w_{2}|\leq C^{\star}\|w_{1}-w_{2}\|^{2}_{\operatorname{L}_{2}(\Omega)}$
@@ 3.5. Lemma (energy estimates for the order parameter)
$\frac{\tau}{2}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\!\!\!$ $+$ $\!\!\!\frac{1}{2}d_{\tau}\|u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+c_{A}|u^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ {\left|\left(g^{\prime}(u^{k}),u^{k}\right)\right|}$ (3.27)
$+{\left|\left((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),u^{k}\right)\right|}+{\left|\left((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),u^{k}\right)\right|}$
$\qquad\qquad\qquad\qquad\qquad\qquad\leq C^{\ast}\|u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)},$
where $C^{\ast}$ is independent of $\tau$. Multiplying by $\tau$ and summing from $k=1,2,...,l$, yields for any $l\leq K$ that
@@
Owing to (3.3) we have $\tau<1/(4C^{\star})$ and hence using a generalized discretized Gronwall lemma, see (Bartels, 2015, Lem 2.2), we have that
@@
which yields the first two bounds in (3.25). To prove the third bound we set $\chi=d_{\tau}u^{k}$ in (3.14), multiply the resulting equation by $\tau$ and sum over $k=1,2,...,l$, to obtain for any $l\leq K$ that
$\tau\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\sum_{k=1}^{l}(\boldsymbol{{A}}(\nabla u^{k})\nabla u^{k},\nabla[u^{k}-u^{k-1}])+\tau\sum_{k=1}^{l}(g^{\prime}(u^{k}),d_{\tau}u^{k})\\ \qquad=\tau\sum_{k=1}^{l}((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),d_{\tau}u^{k})+\tau\sum_{k=1}^{l}((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),d_{\tau}u^{k}).$ (3.30)
@@ 3.6. Lemma (existence and uniqueness of the time-discrete potential)
$b_{1}(u^{k};\phi^{k},\eta)=-(d_{\tau}u^{k}+\sigma^{\prime}(u^{k})\partial_{x}u^{k},\eta).$ (3.34)
@@
$\|d_{\tau}u^{k}+\sigma^{\prime}(u^{k})\partial_{x}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq\frac{C}{\tau}.$ (3.38)
@@ 3.8. Lemma (energy stability of the potential)
$\int_{\Omega}\sigma(u^{k})|\nabla\phi^{k}|^{2}=-\int_{\Omega}d_{\tau}u^{k}\phi^{k}+\int_{\Omega}\sigma^{\prime}(u^{k})\partial_{x}u^{k}\phi^{k}.$
@@
$\leavevmode\nobreak\ \sigma_{\min}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\phi^{k}\|_{\operatorname{L}_{2}(\Omega)}+\sigma^{\prime}_{\max}\|\partial_{x}u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\phi^{k}\|_{\operatorname{L}_{2}(\Omega)},$ (3.46)
@@
$\begin{split}\frac{\sigma_{\min}}{2C_{f}}\|\phi^{k}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\frac{\sigma_{\min}}{2}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq C\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\frac{\sigma_{\min}}{8C_{f}}\|\phi^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\\ +C\|\partial_{x}u^{k}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\frac{\sigma_{\min}}{8C_{f}}\|\phi^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\end{split}$ (3.47)
where $C_{f}$ is the constant in the Friedrichs inequality, which is associated with the diameter of $\Omega$. After rearranging (3.47), multiplying the resulting inequality by $\tau$ and summing over $k=1,2,...,l$, for any $l\leq K$ we have
$\tau\sum_{k=1}^{l}\|\phi^{k}\|^{2}_{\operatorname{H}^{1}(\Omega)}\leq C\Big{(}\tau\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\sum_{k=1}^{l}\|\partial_{x}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\Big{)}$ (3.48)
and the result follows from Lemma 3.5. . ∎

### v3

@@ 3.3. Lemma (stability of the time-discrete order parameter)
$(u^{k},(u^{k})^{-})+\tau(\boldsymbol{{A}}(\nabla u^{k})\nabla u^{k},\nabla(u^{k})^{-})-\tau((m(c^{k-1}))^{-}{\textsl{H}}_{1}(u^{k},u^{k-1}),(u^{k})^{-})\\ -\tau((m(c^{k-1}))^{+}{\textsl{H}}_{2}(u^{k},u^{k-1}),(u^{k})^{-})+\tau(g^{\prime}(u^{k}),(u^{k})^{-})=(u^{k-1},(u^{k})^{-})$ (3.9)
@@
$\|(u^{k})^{-}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\tau c_{A}|(u^{k})^{-}|^{2}_{\operatorname{H}^{1}(\Omega)}\\ \leq(u^{k-1},(u^{k})^{-})+\tau((m(c^{k-1}))^{-}{\textsl{H}}_{1}(u^{k},u^{k-1}),(u^{k})^{-})\\ +\tau((m(c^{k-1}))^{+}{\textsl{H}}_{2}(u^{k},u^{k-1}),(u^{k})^{-})-\tau(g^{\prime}(u^{k}),(u^{k})^{-}).$ (3.10)
By assumption $(u^{k-1})^{-}=0$ and noting from (2.12) and (3.2) that $(g^{\prime}(u^{k}),(u^{k})^{-})\geq 0$, $((m(c^{k-1}))^{-}{\textsl{H}}_{1}(u^{k},u^{k-1}),(u^{k})^{-})\leq 0$ and $((m(c^{k-1}))^{+}{\textsl{H}}_{2}(u^{k},u^{k-1}),(u^{k})^{-})\leq 0$ it follows that
$\|(u^{k})^{-}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq 0\penalty 10000\ \penalty 10000\ \Rightarrow\penalty 10000\ \penalty 10000\ u^{k}\geq 0.$ (3.11)
@@ 3.4. Remark (useful inequalities)
${\left|\left((m(c^{k-1}))^{-}{\textsl{H}}_{1}(u^{k},u^{k-1})+(m(c^{k-1}))^{+}{\textsl{H}}_{2}(u^{k},u^{k-1}),\chi\right)\right|}\leq C\|u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\chi\|_{\operatorname{L}_{2}(\Omega)}.$ (3.16)
3.5. Lemma (existence and uniqueness the time-discrete order parameter)
@@
$(d_{\tau}u^{k},\chi)+I^{\prime}(u^{k})\chi=0,\penalty 10000\ \penalty 10000\ \forall\chi\in\operatorname{H}^{1}(\Omega),$ (3.17)
@@
$\begin{split}I^{\prime}(w)\chi&=\left(\boldsymbol{{A}}(\nabla w)\nabla w,\nabla\chi\right)+\left(g^{\prime}(w)-(m(c^{k-1}))^{-}{\textsl{H}}_{1}(w,u^{k-1}),\chi\right)\\ &\phantom{=}-\left((m(c^{k-1}))^{+}{\textsl{H}}_{2}(w,u^{k-1}),\chi\right)\end{split}$ (3.18)
is the (Fréchet) derivative of the energy functional $I:\operatorname{H}^{1}(\Omega)\to\mathbb{R}$ with, recalling (3.4),
$\begin{split}I(w)&=\int_{\Omega}\Big(\frac{a(\nabla w)^{2}}{2}|\nabla w|^{2}+g(w)-(m(c^{k-1}))^{-}h_{1}(w,u^{k-1})\\ &\phantom{=}-(m(c^{k-1}))^{+}h_{2}(w,u^{k-1})\Big).\end{split}$ (3.19)
@@
$I_{k}(w)=\frac{1}{2\tau}\|w-u^{k-1}\|^{2}_{\operatorname{L}_{2}(\Omega)}+I(w)\ =\int_{\Omega}F(x,w,\nabla w),$ (3.20)
where
$F(x,\xi,\boldsymbol{p}):=\frac{a(\boldsymbol{p})^{2}}{2}|\boldsymbol{p}|^{2}+\frac{1}{2\tau}(\xi-u^{k-1})^{2}\\ +g(\xi)-(m(c^{k-1}))^{-}h_{1}(\xi,u^{k-1})-(m(c^{k-1}))^{+}h_{2}(\xi,u^{k-1}).$ (3.21)
Note that $I_{k}$ is the functional of which (3.17) is the Euler–Lagrange equation. From Theorem 3.30 in Dacorogna (2008), in order to show the existence of a minimiser of (3.20) it is sufficient to exhibit that the function $F$ is a Carathéodory function, coercive and satisfies a suitable growth condition. From (Bartels, 2015, Remark 2.4(i)) it follows that $F$ is a Carathédory function. From (3.14) we deduce that there are constants $\gamma_{1},\gamma_{2}>0$ such that
$\gamma_{1}\leq g(\xi)-(m(c^{k-1}))^{-}h_{1}(\xi,u^{k-1})-(m(c^{k-1}))^{+}h_{2}(\xi,u^{k-1})\leq\gamma_{2}.$ (3.22)
@@
$F(x,\xi,\boldsymbol{p})$ $=\frac{a(\boldsymbol{p})^{2}}{2}|\boldsymbol{p}|^{2}+\frac{1}{2\tau}(\xi-u^{k-1})^{2}+g(\xi)-(m(c^{k-1}))^{-}h_{1}(\xi,u^{k-1})$
$\qquad-(m(c^{k-1}))^{+}h_{2}(\xi,u^{k-1})$
@@
$F(x,\xi,\boldsymbol{p})$ $=\frac{a(\boldsymbol{p})^{2}}{2}|\boldsymbol{p}|^{2}+\frac{1}{2\tau}(\xi-u^{k-1})^{2}+g(\xi)-(m(c^{k-1}))^{-}h_{1}(\xi,u^{k-1})$
$\qquad-(m(c^{k-1}))^{+}h_{2}(\xi,u^{k-1})$
@@
$\int_{\Omega}\frac{1}{\tau}(w_{1}-w_{2})v+\int_{\Omega}(\boldsymbol{{A}}(\nabla w_{1})\nabla w_{1}-\boldsymbol{{A}}(\nabla w_{2})\nabla w_{2})\cdot\nabla v\\ +\,\int_{\Omega}\Big(g^{\prime}(w_{1})-g^{\prime}(w_{2})\Big)v-\int_{\Omega}(m(c^{k-1}))^{+}\Big({\textsl{H}}_{1}(w_{1},u^{k-1})-{\textsl{H}}_{1}(w_{2},u^{k-1})\Big)v\\ -\int_{\Omega}(m(c^{k-1}))^{-}\Big({\textsl{H}}_{2}(w_{1},u^{k-1})-{\textsl{H}}_{2}(w_{2},u^{k-1})\Big)v=0.$ (3.25)
Choosing $v=w_{1}-w_{2}$ in (3.25), using the Lipschitz continuity of $g^{\prime}$, ${\textsl{H}}_{1}(u^{k},u^{k-1})$ and ${\textsl{H}}_{2}(u^{k},u^{k-1})$, (3.14), and noting from Remark 2.6 the monotonicity of the mapping $w\mapsto J_{a}^{\prime}(w)$, we obtain
$\frac{1}{\tau}\|w_{1}-w_{2}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq\int_{\Omega}|(g^{\prime}(w_{1})-g^{\prime}(w_{2}))(w_{1}-w_{2})|\\ +C\int_{\Omega}|{\textsl{H}}_{1}(w_{1},u^{k-1})-{\textsl{H}}_{1}(w_{2},u^{k-1})||w_{1}-w_{2}|\\ \leq C^{\star}\|w_{1}-w_{2}\|^{2}_{\operatorname{L}_{2}(\Omega)}$ (3.26)
@@ 3.6. Lemma (energy estimates for the order parameter)
$\frac{\tau}{2}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{1}{2}d_{\tau}\|u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+c_{A}|u^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\\ \leq{\left|\left(g^{\prime}(u^{k}),u^{k}\right)\right|}\\ +{\left|\left((m(c^{k-1}))^{-}{\textsl{H}}_{1}(u^{k},u^{k-1}),u^{k}\right)\right|}+{\left|\left((m(c^{k-1}))^{+}{\textsl{H}}_{2}(u^{k},u^{k-1}),u^{k}\right)\right|}\\ \qquad\qquad\qquad\qquad\qquad\qquad\leq C^{\ast}\|u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)},$ (3.30)
where $C^{\ast}$ is independent of $\tau$. Multiplying by $\tau$ and summing from k=1,2,,lk=1,2,...,l, yields for any $l\leq K$ that
@@
Owing to (3.6) we have $\tau<1/(4C^{\star})$. Hence using a generalized discretized Gronwall lemma (Bartels, 2015, Lem 2.2) we obtain
@@
which yields the first two bounds in (3.28). To prove the third bound we set $\chi=d_{\tau}u^{k}$ in (3.17), multiply the resulting equation by $\tau$ and sum over k=1,2,,lk=1,2,...,l, to obtain for any $l\leq K$ that
$\tau\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\sum_{k=1}^{l}(\boldsymbol{{A}}(\nabla u^{k})\nabla u^{k},\nabla[u^{k}-u^{k-1}])+\tau\sum_{k=1}^{l}(g^{\prime}(u^{k}),d_{\tau}u^{k})\\ \qquad=\tau\sum_{k=1}^{l}((m(c^{k-1}))^{-}{\textsl{H}}_{1}(u^{k},u^{k-1}),d_{\tau}u^{k})+\tau\sum_{k=1}^{l}((m(c^{k-1}))^{+}{\textsl{H}}_{2}(u^{k},u^{k-1}),d_{\tau}u^{k}).$ (3.33)
@@ 3.7. Lemma (existence and uniqueness of the time-discrete potential)
$b_{1}(u^{k};\phi^{k},\eta)=-(d_{\tau}u^{k}+\sigma^{\prime}(u^{k})\partial_{x_{1}}u^{k},\eta).$ (3.37)
@@
$\|d_{\tau}u^{k}+\sigma^{\prime}(u^{k})\partial_{x_{1}}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq\frac{C}{\tau}.$ (3.41)
@@ 3.9. Lemma (energy stability of the potential)
$\int_{\Omega}\sigma(u^{k})|\nabla\phi^{k}|^{2}=-\int_{\Omega}d_{\tau}u^{k}\phi^{k}+\int_{\Omega}\sigma^{\prime}(u^{k})\partial_{x_{1}}u^{k}\phi^{k}.$
@@
$\penalty 10000\ \sigma_{\min}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\phi^{k}\|_{\operatorname{L}_{2}(\Omega)}+\sigma^{\prime}_{\max}\|\partial_{x_{1}}u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\phi^{k}\|_{\operatorname{L}_{2}(\Omega)},$ (3.49)
@@
$\begin{split}\frac{\sigma_{\min}}{2C_{f}}\|\phi^{k}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\frac{\sigma_{\min}}{2}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq C\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\frac{\sigma_{\min}}{8C_{f}}\|\phi^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\\ +C\|\partial_{x_{1}}u^{k}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\frac{\sigma_{\min}}{8C_{f}}\|\phi^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\end{split}$ (3.50)
where $C_{f}$ is the constant in the Friedrichs inequality, which is associated with the diameter of $\Omega$. After rearranging (3.50), multiplying the resulting inequality by $\tau$ and summing over k=1,2,,lk=1,2,...,l, for any $l\leq K$ we have
[…]
and the result follows from Lemma 3.6. ∎